OpenStudy (anonymous):

An artist is designing a kite like the one show below. Calculate the area to determine how much material she will need to create the kite. A kite with two top triangles with height of 7 inches and a base 9.5 inches. The bottom two triangles share the base but have a height labeled 12 inches. 38 square inches 66.5 square inches 114 square inches 180.5 square inches @PlanSlam @PaxPolaris

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

so it will be C

OpenStudy (anonymous):

thanks! @PlanSlam

OpenStudy (anonymous):

np bro

OpenStudy (paxpolaris):



OpenStudy (paxpolaris):

 area of half the kite \[=\frac12 bh\\=\frac12\left( 12+7 \right)\times 9.5\\=\frac12 13 \times 9.5\]

OpenStudy (paxpolaris):

sorry \[=\frac12 \times 19 \times 9.5\]

OpenStudy (paxpolaris):

so the are of the entire kite will be twice that 19*9.5 = 180.5 D is the correct answer @sjman1

OpenStudy (anonymous):

@PaxPolaris

OpenStudy (anonymous):

is it still right @PaxPolaris

OpenStudy (paxpolaris):

that is the shape i drew . D is correct